OpenStudy (anonymous):

Use the substitution method to find the solutions to the systems of equations Y=-x^2 Y=x Please help got to do today

OpenStudy (anonymous):

x= o and -1

OpenStudy (anonymous):

How did you get that

OpenStudy (anonymous):

we have y=x=-x^2

OpenStudy (anonymous):

or simply x=-x^2

OpenStudy (anonymous):

shift everything to LHS, so x^2 +x=0

OpenStudy (anonymous):

take x common so the equation is x(x+1)=0 now the product of two numbers( x nd (x+1)) is 0, so either one of them can be equal to 0

OpenStudy (anonymous):

so this implies either x=0 or x+1=0 solving both we get x= 0 and -1